A tubular steel strut of outer diameter 20 mm and inner diameter 5 mm carries a compressive load of 10 kN. The strut is 1 m long and the modulus of elasticity of the material is 60 GPa. Calculate the induced compressive stress, compressive strain and the change in length that takes place during the loading.
1
Expert's answer
2020-06-02T17:54:28-0400

Cross-sectional area of the strut


"A=0.25\pi (0.02^2-0.005^2)=0.00029\ m^2"

Compressive Stress, 


"\sigma=\frac{10\cdot10^3}{0.00029}=34.5\ Mpa"

Compressive Strain 


"ϵ=\frac{34.5\cdot10^6}{60\cdot10^9}=5.75\cdot10^{-4}"

"Δ
L
=
ϵ

L
\\=5.75\cdot10^{-4}(1)=5.75\cdot10^{-4}m=0.575\ mm"